Define (a) electric intensity (b) difference of a potential. How are these quantities related?
1
Expert's answer
2020-11-06T10:18:16-0500

The electric current force is the value "I" , which characterizes the ordered movement of electric charges and which is numerically equal to the amount of charge "\u0394q" flowing through a given surface "S" (cross section of the conductor) per unit time: "I =\frac {Δq} {t}" .

Electric voltage (potential difference) between points A and B of an electric circuit or electric field is a scalar physical value, the value of which is equal to the operation of an effective electric field (including third-party fields), performed when transferring a single test electric charge from point A to point B.

"U=-\Delta\phi=\frac{A}{q}" .

The Ohm law is an empirical physical law that determines the connection of electric voltage with the current intensity flowing in the conductor and the resistance of the conductor. "R=\frac{U}{I}"